Translate this to a mathematical equation and solve .
(please show work)
In four rounds of a card game, you get scores of -3, 9, 8 and -3. What is the final score ??

Answer by karaoz(32)


What is a card game? Assuming some simple additive card games, then
The final score = score in game 1 + score in game 2 + score in game 3 + score in game 4.
If we let the final score be F and scores in game n to be Gn then we have a simpler equations (at least in terms of its length):
F = G1 + G2+ G3 + G4
and since we know G1, G2, G3 and G4, we can write:
F = -3 + 9 + 8 +(-3),
which is
F = -3 + 9 + 8 - 3,
which is further equivalent to:
F = 11.

11 is a good number, which is the only good thing about this question.